Find the composite function f1 with respect to f2 , such that f1={ (2,3),(4,9),(8,1)} and f2={ (5,2),(3,4),(7,8)}.
Here is my answer: (5,3), (3,9), (7,1)
Am I correct? Please show the working, or give an explanation.
Answer by rothauserc(4718) 
 you are correct
we are asked to find f1 o f2(x) where x is an element of the domain of f2
note that the given points are of the form (x, f(x))
f1 o f2(5) = f1(2) = 3
f1 o f2(3) = f1(4) = 9
f1 o f2(7) = f1(8) = 1
therefore,
f1 o f2(x) = (5,3), (3,9), (7,1)
|
|
|